# Anomalous radiation reaction in a circularly polarized field

### v6

@@ III Results and Discussion
$q_{m}=\frac{1}{\lambda_{0}}\left(\sqrt{[1-\left(\mathbf{v}\cdot\mathbf{n}^{(\mathbf{q})}\right)/c]^{2}+2m\lambda_{0}\omega/c}-[1-(\mathbf{v}\cdot\mathbf{n}^{(\mathbf{q})})/c]\right)=\frac{m\omega}{c}\left[1+\frac{\left(\mathbf{v}\cdot\mathbf{n}^{(\mathbf{q})}\right)}{c}+{\it O}(1/c^{2})\right],$ (A15)
@@
where the symbol $\cal{P}$} denotes the principal value of the integral (see, e.g., Sec. 43 in Ref. Landau_3, ). Calculating the first term in the right-hand side of Eq. (A28), one can see that it makes only the relativistically small contribution to the classical recoil force (A22) and, therefore, will be neglected in the following analysis. As to the second delta-function term accounting for the on-shell photons, it yields the substantially new addition ${w}_{\mathbf{k}}(\mathbf{q})\propto v_{{\mathbf{k}}y,x}q_{x,y}$ to the probability (III), which crucially modifies the electron dynamics discussed below. Keeping in mind that the field (A1) is assumed to be strong ($v_{0}\gg v_{\mathbf{k}}$), the total probability amplitude, $T_{\mathbf{k}}(\mathbf{q})=T^{(a)}_{\mathbf{k}}(\mathbf{q})+T^{(b)}_{\mathbf{k}}(\mathbf{q})$, can be written in the leading order of expansion in powers of $1/c$ as

### v7

@@ III Results and Discussion
$q_{m}=\frac{1}{\lambda_{0}}\left(\sqrt{[1-\left(\mathbf{v_{k}}\cdot\mathbf{n}^{(\mathbf{q})}\right)/c]^{2}+2m\lambda_{0}\omega/c}-[1-(\mathbf{v_{k}}\cdot\mathbf{n}^{(\mathbf{q})})/c]\right)=\frac{m\omega}{c}\left[1+\frac{\left(\mathbf{v_{k}}\cdot\mathbf{n}^{(\mathbf{q})}\right)}{c}+{\it O}(1/c^{2})\right],$ (A15)
@@
where the symbol $\cal{P}$} denotes the principal value of the integral (see, e.g., Sec. 43 in Ref. Landau_3, ). Calculating the first term in the right-hand side of Eq. (A28), one can see that it makes only the relativistically small contribution to the classical recoil force (A22) and, therefore, will be neglected in the following analysis. As to the second delta-function term accounting for the on-shell photons, it yields the substantially new addition ${w}_{\mathbf{k}}(\mathbf{q})\propto v_{{\mathbf{k}}y,x}q_{x,y}$ to the probability (III), which crucially modifies the electron dynamics discussed below. Keeping in mind that an electron is assumed to be slow enough ($v_{\mathbf{k}}\ll v_{0}\ll c$), the total probability amplitude, $T_{\mathbf{k}}(\mathbf{q})=T^{(a)}_{\mathbf{k}}(\mathbf{q})+T^{(b)}_{\mathbf{k}}(\mathbf{q})$, can be written in the leading order of expansion in powers of $1/c$ as